Modern electroplaters can cover a surface area of 60.0 m2 with one troy ounce of gold (volume 1.611 cm3). What is the thickness of the electroplated gold? 2.68 10-8 m 1.34 10-9 m 1.67 10-6 m 3.33 10-7 m
Answer #1

2 Area Volume = 1-6|1Crn3 v= 1-6/1x106m 268 o